find the force of gravitation between two objects weighing 450kg and 570kg when they are 0.750 m apart
1
Expert's answer
2014-08-05T11:11:42-0400
If one can assume objects as material points,
F = G m1 m2 /r2=6.67*10^-11*450*570/0.75^2=3*10^-5 N
otherwise force will be strongly dependent on shape and mass distribution inside objects.